VIBRATIONAL SPECTROSCOPY OF GERMANIUM-CARBON CLUSTERS: ν_{4}(σ_{u}) MODE OF GeC_{5}Ge

\underline{ERIC GONZALEZ}, C. M. L. RITTBY, and W. R. M. GRAHAM, Department of Physics and Astronomy, Texas Christian University,Fort Worth, TX 76129.

Recent results will be presented from FTIR (Fourier transform infrared) and DFT (density functional theory) studies of the vibrational fundamentals and structures of germanium-carbon clusters trapped in solid Ar, which have previously included linear GeC_{3}Ge^{a} and GeC_{3}Si^{b}. The linear germanium-carbon cluster GeC_{5}Ge has been detected using FTIR spectra generated when products from the dual laser evaporation of Ge and C rods are trapped in solid Ar matrices at ∼10 K. Comparison of frequencies and ^{13}C isotopic shift measurements with the predictions of DFT calculations at the B3LYP/cc-pVDZ level confirms the identification of the ν_{4}(σ_{u}) mode of GeC_{5}Ge at 2158.0 cm^{−1}.

aD. L. Robbins, C. -M. L Rittby, W. R. M Graham,J. Chem. Phys. \underline{114}, 3570 (2000).

bD. L. Robbins, C. M. L Rittby, W. R. M Graham,J. Chem. Phys. \underline{117}, 3811 (2002).